What is the vapor pressure of the pure solvent if the vapor pressure of a solution of 15g sucrose (C12H22O11) in 100g of ethanol (C2H5OH) is 58 mmHg?

1
Expert's answer
2022-02-17T13:13:02-0500

Molar mass of sucrose = 342g/mol

Moles of sucrose = "\frac{15}{342}=0.04386"

Molar mass of ethanol= 46g/mol

Moles of ethanol= 100/46=2.1739

Total moles = 0.04386+2.1739=2.21776

Mole fraction of ethanol = 2.1739/2.21776=0.98022

"58=0.98022\u00d7P_{ethanol}"

"P \nethanol\n\u200b\n =59.17mmHg"